OpenStudy (anonymous):

solve the inequality 7x+5 (is less than equal) to 19

OpenStudy (anonymous):

\[7x+5\le19\]

OpenStudy (anonymous):

\[\Large 7x\leq 19-5\] divide by 7 both parts and there you go !

OpenStudy (anonymous):

\[ 7 x + 5 \le 19\\ 7 x \le 14\\ x \le 2 \]